Please answer the following cash flow diagrams by finding the missing value. Note that downward arrows represent deposits and upward arrows represent withdrawals.

/Mo مو : 42 (f) [2 points) 2 ہ 3 4 . 5 و (g) 3 points] | / +57/v هم موير = 72 5 ، 2 في 4 ک ۶۰۰۰ ه -۷ و
Answer #1

Solution :- (1)

B | 1 A Solution - 1 Years Cashflows PVE @ 6% 0.943 -800 0.890 -800 0.840 0.792 200 0.747 350 0.705 500 0.665 650 0.627 0.592

1 Solution :-1 Years Cashflows -800 -800 3 4 5 1 6 2 73 1 8 4 95 106 117 128 PVF @ 6% =1/1.06 =C5/1.06 =C6/1.06 =C7/1.06 =C8/

The Negative Amount shows the Deposits and the Positive amount shows the withdrawls

Now P is the Present Value of Deposits and Withdrawls which is equal to $226.36

Here in the figure we see the amount after 5th year regularly increase by 150 so we add 150

And if there is any doubt please ask in comments

As per HomeworkLib policy we need to answer only one question at once so please ask others as seperate one